Dennis Co. shows the following information on its 2016 income statement: sales = $157,000; costs = $81,100; other expenses = $4,400; depreciation expense = $10,100; interest expense = $7,600; taxes = $18,830; dividends = $7,600. In addition, you're told that the firm issued $3,400 in new equity during 2016 and redeemed $5,500 in outstanding long-term debt.
  
a. What is the 2016 operating cash flow? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
  
Operating cash flow            $
  
b. What is the 2016 cash flow to creditors? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
  
Cash flow to creditors            $
  
c. What is the 2016 cash flow to stockholders? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
  
Cash flow to stockholders            $
  
d. If net fixed assets increased by $20,650 during the year, what was the addition to NWC? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
  
Addition to net working capital            $
Answer #1

A: OCF = Sales – costs- Other expenses -Taxes

= 157000- 81100- 4400- 18830

=52670

B: Cash flow to creditors = Interest expenses + Debt redeemed

= 7600 + 5500

=13100

C: Cash flow to stockholders = Dividend- New equity

= 7600-3400

= $4200

D: Cash flow from assets = OCF – Capital spending – Additions to NWC

Cash flow to creditors + Stockholders = OCF – Capital spending – Additions to NWC

13100+ 4200 = 52670- 20650- Additions to NWC

Additions to NWC= 14720